OpenStudy (anonymous):

x^2+20x=30 i need this square completed and its * two answers (*if answer is not an integer round to nearest tenth . thank you .!

OpenStudy (anonymous):

idont understand the way the answers on that site are written .

OpenStudy (nikita2):

Oh, then use the formula: x1/x2 = -10 +/- sqrt(D), where D = (b/2)^2 - ab, where b = 20, a = 1, c = -30

OpenStudy (anonymous):

uhhhh im so lost i hate math